# Matérn Kernels for Tunable Implicit Surface Reconstruction

### v2

@@ 1 Introduction
where the function space $H:=H(\Omega)$ is usually taken to be a Reproducing Kernel Hilbert Space (RKHS) with associated reproducing kernel $k:\Omega\times\Omega\longrightarrow\mathbb{R}$, and $\lambda>0$. Once $\hat{f}$ at hand, the reconstructed object’s surface is implicitly given as $\mathcal{\hat{S}}=\{x:\hat{f}(x)=0\}\subset\mathbb{R}^{d}$ and can be extracted using Marching Cubes (Lorensen & Cline, 1987). It is quite easy to observe that, as $\lambda\rightarrow 0$, the optimization problem in Eq. (1) turns into the constrained minimization problem,
@@
which makes it easier to see how the predicted surface behaves away from the input points. As seen from Eq. (2), while $\mathcal{\hat{S}}$ should interpolate the given points $\mathcal{X}$ exactly, the behavior in between and away from those points is solely determined by the induced norm $\|f\|_{H}$ of the function space, $H$, over which we are optimizing. The properties of $\|f\|_{H}$ are uniquely defined by the reproducing kernel, $k$; this is a direct consequence of the Representer Theorem (Kimeldorf & Wahba, 1970; Schölkopf et al., 2001) which states that solutions to Eq. (1) are of the form $f(x)=\sum_{i=1}^{m}\alpha_{i}k(x,x_{i})$, yielding $\|f\|_{H}^{2}=\alpha^{\top}K\alpha$. Here, $\alpha\in\mathbb{R}^{m}$ and $K=(K)_{ij}=k(x_{i},x_{j})\in\mathbb{R}^{m\times m}$. This shows that the behavior of the norm and hence, estimated surfaces $\mathcal{\hat{S}}$, can be explicitly controlled by the kernel function, enabling easy injection of inductive biases (such as smoothness assumptions) into the surface reconstruction problem.
@@
Although already used in the mid-90s and early 2000s (Savchenko et al., 1995; Carr et al., 1997; 2001), only recently, kernel-based methods for 3D implicit surface reconstruction became extremely competitive, with Neural Kernel Surface Reconstruction (NKSR; Huang et al. (2023)) eventually evolving as the new state-of-the-art. While early works mostly focused on (polyharmonic) radial basis functions such as thin-plate splines, recent work (Williams et al., 2021; 2022) employ the first-order arc-cosine kernel. Introduced by Cho & Saul (2009), arc-cosine kernels have been shown to mimic the computation of (two-layer) fully-connected ReLU networks; indeed, if the layer-width tends to infinity, the first-order arc-cosine kernel is identified with the Neural Tangent Kernel (NTK; Jacot et al. (2018)) of the network. One important aspect of this kernel is that it does not have adjustable parameters. While this can be advantageous in some situations, it also prevents users from tuning their reconstructions. If the arc-cosine kernel fails to accurately recover a surface, then there is no possibility (except for using more observations) to improve the result.
@@
In this work, we suggest a different family of kernels (with parameters $\nu,h>0$) for implicit surface reconstruction: Matérn kernels (Matérn, 1986; Stein, 1999), see Figures 1 and 2. Contrary to arc-cosine kernels, Matérn kernels are stationary (hence translation invariant), spatially decaying (thus leading to sparse Gram matrices), and unify a variety of popular kernel functions, including the Laplace and Gaussian kernel. As we will show, Matérn kernels have appealing properties, making them the ideal candidate for kernel-based surface reconstruction. From a practical perspective, we demonstrate that a simple change of the kernel function from arc-cosine to Matérn leads to a consistently improved reconstruction accuracy and a significant speedup. From a theoretical perspective, we argue that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings (Tancik et al., 2020) enable coordinate-based Multi-layer Perceptrons (MLPs) to learn high-frequency functions in low-dimensional domains, effectively overcoming spectral bias (the slow convergence towards high-frequency details). In addition, we prove that, as the layer-width approaches infinity, Matérn kernels are identified with the NTK of two-layer Siren (fully-connected MLPs with sine activations; Sitzmann et al. (2020)) networks—together with Fourier feature mappings the two arguably most influential methods to overcome spectral bias in coordinate-based MLPs. Lastly, we establish a connection between arc-cosine and Matérn kernels by showing that the Laplace kernel (Matérn kernel with $\nu=1/2$) and the arc-cosine kernel are equal up to an affine transformation when restricted to the sphere. In summary, our key contributions are:
@@
  • We propose data-dependent (i.e., learnable) Matérn kernels by leveraging the Neural Kernel Field (NKF) framework (Williams et al., 2022).
@@ 2 Related Work
We briefly review some relevant literature about 3D implicit surface reconstruction from oriented point clouds, focusing on kernel-based methods. For a more in-depth overview, including neural-network-based reconstruction methods, see surveys by Berger et al. (2017); Huang et al. (2022b).
Early kernel-based 3D surface reconstruction methods mostly employ Radial Basis Functions (RBFs) such as thin-plate splines (Savchenko et al., 1995), biharmonic RBFs (Carr et al., 1997; 2001), or Gaussian kernels (Schölkopf et al., 2005). Nowadays, the most widely used surface reconstruction technique is Screened Poisson Surface Reconstruction (SPSR; Kazhdan & Hoppe (2013)) which, however, can itself be viewed as kernel method (Williams et al., 2021). Only recently, Neural Splines (NS; Williams et al. (2021)) proposed to use the so-called (first-order) arc-cosine kernel,
@@
for implicit surface reconstruction, achieving state-of-the-art results that outperform classical surface reconstruction techniques and non-linear methods based on neural networks by a large margin. This method laid the cornerstone for Neural Kernel Fields (NKFs; Williams et al. (2022)) which attempts to make arc-cosine kernels learnable by passing input points through a task-specific neural network before evaluating the kernel function, similar to Deep Kernel Learning (Wilson et al., 2016). Based on SPSR, NeuralGalerkin (Huang et al., 2022a) proposed to learn basis functions inferred by a sparse convolutional network instead of using a fixed Bézier basis as in SPSR, hence can be seen as kernel method in the broader sense. Neural Kernel Surface Reconstruction (NKSR; Huang et al. (2023)) built upon NKF and NeuralGalerkin and proposed an all-purpose, highly scalable surface reconstruction method that is robust against noise and eventually became state-of-the-art.
@@ 3 Matérn Kernels for Tunable Surface Reconstruction
We propose to use the family of Matérn kernels (Matérn, 1986; Stein, 1999) for implicit surface reconstruction, being defined as
@@
For more information about Matérn kernels, please refer to Porcu et al. (2023). Next, we revisit some basic properties of Matérn kernels and place them in the context of surface reconstruction.
@@ 3.1 Basic Properties
Stationary. Matérn kernels are stationary, i.e., they only depend on the difference $x-y$ between two points $x,y\in\Omega$. If the distance is Euclidean, Matérn kernels are also isotropic; they only depend on $\|x-y\|$. As a result, Matérn kernels are rotation and translation invariant, hence being independent of the absolute embedding of points $\mathcal{X}$ (based on the reproducing property, translating input points does not change the reconstructed surface). Consequently, since the kernel’s value only depends on the relative positions, objects in a scene (or parts of an object) with similar geometric properties will be reconstructed consistently as long as the relative distances between points on each instance (or part) remain the same. This is in contrast to the arc-cosine kernel, whose value depends on the absolute position of points, see Eq. (3); hence, it is not translation invariant (thus non-stationary), and multiple instances of an object in a scene will generally not yield the same surface reconstruction. See Appendix A for further discussion.
Spatially decaying. As opposed to arc-cosine kernels, Matérn kernels are spatially decaying; they tend to zero as the distance $\|x-y\|$ becomes large. Consequently, although Matérn kernels are technically not compact (or locally supported), because their value decays exponentially fast (Porcu et al., 2023), they can be considered effectively compact. From a computational perspective, this is a very desirable property as it leads to sparse Gram matrices when truncated with a special kernel (Genton, 2001), allowing the use of highly efficient and scalable sparse linear solvers. Finally, we note that simply truncating any kernel does not yield a valid, positive definite kernel in general.
@@ 3.2 Eigenvalue Decay, Spectral Bias, and Reconstruction Error
Eigenvalue decay and spectral bias. We begin by showing that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings help coordinate-based MLPs overcome spectral bias. As investigated in Tancik et al. (2020), a rapid decrease of the NTK’s eigenvalues implies an associated MLP’s slow convergence to high-frequency components of the target function (up to the point where the network is practically unable to learn these components). Consequently, a smaller eigenvalue decay rate (EDR), i.e., a slower eigenvalue decay, leads to faster convergence to high-frequency content—in the context of implicit surface reconstruction, more detailed geometry. To overcome this slow convergence, a phenomenon known as spectral bias, Tancik et al. (2020) use a Fourier feature mapping of the form
@@
Theorem 1 (Seeger et al. (2008), Theorem 3).
@@
Figure 3 visualizes the EDR for different values of $\nu$ and $h$ along with the EDR of the non-tunable, first-order arc-cosine kernel which equates to $\Theta(s^{-(1+d)/d})$ and has only been very recently proven by Li et al. (2024) for general input domains and distributions. As seen, the smoother the kernel (i.e., the larger $\nu$) and the greater $h$, the faster the eigenvalues decay. This implies less detailed surface reconstruction (with a higher error) for smoother kernels and larger $h$; indeed, this is exactly what we observe in practice, see Figure 4. Contrary, by lowering both parameters, one can achieve faster convergence to high-frequency geometric details, effectively leading to more nuanced surface reconstruction and a smaller reconstruction error.
Figure 4: Matérn kernels are tunable. Surface reconstructions can be improved in practice by varying the bandwidth parameter, $h$, effectively tuning the kernels’ EDR (see also Figure 3). However, setting $h$ too small ($<0.5$) results in overfitting, while setting $h$ too big ($>10$) oversmooths (i.e., underfits) the true surface. This is also reflected in the reconstruction error (measured using Chamfer distance), plotted as a function of $h$ on the right.
@@
is the norm of the RKHS induced by a Matérn kernel with bandwidth $h$ and smoothness $\nu$. Moreover, $C_{d,\nu}:=(2^{d}\pi^{d/2}\Gamma(\nu+d/2)(2\nu)^{\nu})/\Gamma(\nu)$, and $\mathcal{F}[f]$ denotes the Fourier transform of a function $f$. For more information on the bound, including details on $C_{\mathcal{X},\Omega}$, please see, e.g., Santin & Schaback (2016). By inspecting Eq. (10), we notice that the magnitude of the RKHS norm has a considerably high influence on the reconstruction error. Based on this observation and the fact that $\|f\|_{H_{\nu}}$ depends on $h$, our goal is now to further study the effect of $h$ on the RKHS norm. As a first result, we have:
@@ 3.3 Relation to Neural Networks and the Arc-cosine Kernel
Relation to neural networks. Next, we study the connection between Matérn kernels and neural networks. It is well known that the first-order arc-cosine kernel mimics the computation of a two-layer, infinite-width ReLU network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$, $f(x)=m^{-1/2}\sum_{i=1}^{m}v_{i}[w^{\top}_{i}x+b_{i}]_{+}$ when the bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$ are fixed from initialization and drawn from a standard normal distribution, see, e.g., Cho & Saul (2009); Williams et al. (2021). Here, $[x]_{+}:=\max\{0,x\}$ denotes the ReLU activation function.
Under similar assumptions, we now show that Matérn kernels mimic the computation of two-layer Sirens (Sitzmann et al., 2020) if the layer-width approaches infinity. We claim the following:
@@
Theorem 4 (Chen & Xu (2021), Theorem 1).
@@
When restricted to the hypersphere $\mathbb{S}^{d-1}$, (1) the RKHS of the Matérn kernel with smoothness $\nu=1/2$ (the Laplace kernel) is equal to the RKHS induced by the first-order arc-cosine kernel, implying that (2) the Laplace and arc-cosine kernel are equal up to an affine transformation.
See Appendix E for a proof. For general input domains, our empirical results presented next support (at least for ($L=2$)-layer networks) the widely believed claim that the NTK is not significantly different from standard kernels (Belkin et al., 2018; Hui et al., 2019; Geifman et al., 2020), such as the Laplace kernel. Indeed, we show that the Laplace kernel even outperforms the arc-cosine kernel.

### v3

@@ 1 Introduction
where the function space $H:=H(\Omega)$ is usually taken to be a Reproducing Kernel Hilbert Space (RKHS) with associated reproducing kernel $k:\Omega\times\Omega\longrightarrow\mathbb{R}$, and $\lambda>0$. Once $\hat{f}$ at hand, the reconstructed object’s surface is implicitly given as $\mathcal{\hat{S}}=\{x:\hat{f}(x)=0\}\subset\mathbb{R}^{d}$ and can be extracted using Marching Cubes (lorensen1987). It is quite easy to observe that, as $\lambda\rightarrow 0$, the optimization problem in Eq. (1) turns into the constrained minimization problem,
@@
which makes it easier to see how the predicted surface behaves away from the input points. As seen from Eq. (2), while $\mathcal{\hat{S}}$ should interpolate the given points $\mathcal{X}$ exactly, the behavior in between and away from those points is solely determined by the induced norm $\|f\|_{H}$ of the function space, $H$, over which we are optimizing. The properties of $\|f\|_{H}$ are uniquely defined by the reproducing kernel, $k$; this is a direct consequence of the Representer Theorem (kimeldorf1970; schoelkopf2001) which states that solutions to Eq. (1) are of the form $f(x)=\sum_{i=1}^{m}\alpha_{i}k(x,x_{i})$, yielding $\|f\|_{H}^{2}=\alpha^{\top}K\alpha$. Here, $\alpha\in\mathbb{R}^{m}$ and $K=(K)_{ij}=k(x_{i},x_{j})\in\mathbb{R}^{m\times m}$. This shows that the behavior of the norm and hence, estimated surfaces $\mathcal{\hat{S}}$, can be explicitly controlled by the kernel function, enabling easy injection of inductive biases (such as smoothness assumptions) into the surface reconstruction problem.
@@
Although already used in the mid-90s and early 2000s (savchenko1995; car1997; car2001), only recently, kernel-based methods for 3D implicit surface reconstruction became extremely competitive, with Neural Kernel Surface Reconstruction (NKSR; huang2023) eventually evolving as the new state-of-the-art. While early works mostly focused on (polyharmonic) radial basis functions such as thin-plate splines, recent work (williams2021; williams2022) employ the first-order arc-cosine kernel. Introduced by cho2009, arc-cosine kernels have been shown to mimic the computation of two-layer fully-connected ReLU networks; indeed, in the infinite-width limit and with fixed bottom layer weights, the first-order arc-cosine kernel becomes the Neural Tangent Kernel (NTK; jacot2018) of the network. One important aspect of this kernel is that it does not have adjustable parameters. While this can be advantageous in some situations, it also prevents users from tuning their reconstructions. If the arc-cosine kernel fails to accurately recover a surface, then there is no possibility (except for using more observations) to improve the result.
@@
In this work, we suggest a different family of kernels (with parameters $\nu,h>0$) for implicit surface reconstruction: Matérn kernels (matern1986; stein1999), see Figures 1 and 2. Contrary to arc-cosine kernels, Matérn kernels are stationary (hence translation invariant), spatially decaying (thus leading to sparse Gram matrices), and unify a variety of popular kernel functions, including the Laplace and Gaussian kernel. As we will show, Matérn kernels have appealing properties, making them the ideal candidate for kernel-based surface reconstruction. From a practical perspective, we demonstrate that a simple change of the kernel function from arc-cosine to Matérn leads to a consistently improved reconstruction accuracy and a significant speedup. From a theoretical perspective, we argue that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings (tancik2020) enable coordinate-based Multi-layer Perceptrons (MLPs) to learn high-frequency functions in low-dimensional domains, effectively overcoming spectral bias (the slow convergence towards high-frequency details). In addition, we prove that, as the layer-width approaches infinity, with fixed bottom layer weights and under certain initializations, Matérn kernels are identified with the NTK of two-layer Siren (fully-connected MLPs with sine activations; sitzmann2020) networks—together with Fourier feature mappings the two arguably most influential methods to overcome spectral bias in coordinate-based MLPs. Lastly, we establish a connection between arc-cosine and Matérn kernels by showing that the Laplace (Matérn kernel with $\nu=1/2$) and first-order arc-cosine kernel are equal up to scaling and addition of another kernel function when restricted to the sphere. In summary, our key contributions are:
@@
  • We propose data-dependent (i.e., learnable) Matérn kernels by leveraging the Neural Kernel Field (NKF) framework (williams2022).
@@ 2 Related Work
We briefly review some relevant literature about 3D implicit surface reconstruction from oriented point clouds, focusing on kernel-based methods. For a more in-depth overview, including neural-network-based reconstruction methods, see surveys by berger2017; huang2022.
Early kernel-based 3D surface reconstruction methods mostly employ Radial Basis Functions (RBFs) such as thin-plate splines (savchenko1995), biharmonic RBFs (car1997; car2001), or Gaussian kernels (schoelkopf2005). Nowadays, the most widely used surface reconstruction technique is Screened Poisson Surface Reconstruction (SPSR; kazhdan2013) which, however, can itself be viewed as kernel method (williams2021). Only recently, Neural Splines (NS; williams2021) proposed to use the so-called (first-order) arc-cosine kernel,
@@
for implicit surface reconstruction, achieving state-of-the-art results that outperform classical surface reconstruction techniques and non-linear methods based on neural networks by a large margin. This method laid the cornerstone for Neural Kernel Fields (NKFs; williams2022) which attempts to make arc-cosine kernels learnable by passing input points through a task-specific neural network before evaluating the kernel function, similar to Deep Kernel Learning (wilson2016). Based on SPSR, NeuralGalerkin (huang2022_ng) proposed to learn basis functions inferred by a sparse convolutional network instead of using a fixed Bézier basis as in SPSR, hence can be seen as kernel method in the broader sense. Neural Kernel Surface Reconstruction (NKSR; huang2023) built upon NKF and NeuralGalerkin and proposed an all-purpose, highly scalable surface reconstruction method that is robust against noise and eventually became state-of-the-art.
@@ 3 Matérn Kernels for Tunable Surface Reconstruction
We propose to use the family of Matérn kernels (matern1986; stein1999) for implicit surface reconstruction, being defined as
@@
For more information about Matérn kernels, please refer to porcu2023. Next, we revisit some basic properties of Matérn kernels and place them in the context of surface reconstruction.
@@ 3.1 Basic Properties
Stationary. Matérn kernels are stationary, i.e., they only depend on the difference xyx-y between two points $x,y\in\Omega$. If the distance is Euclidean, Matérn kernels are also isotropic; they only depend on $\|x-y\|$. As a result, Matérn kernels are rotation and translation invariant, hence being independent of the absolute embedding of points $\mathcal{X}$ (based on the reproducing property, translating input points does not change the reconstructed surface). Consequently, since the kernel’s value only depends on the relative positions, objects in a scene (or parts of an object) with similar geometric properties will be reconstructed consistently as long as the relative distances between points on each instance (or part) remain the same. This is in contrast to the arc-cosine kernel, whose value depends on the absolute position of points, see Eq. (3); hence, it is not translation invariant (thus non-stationary), and multiple instances of an object in a scene will generally not yield the same surface reconstruction. See Appendix A for further discussion.
Spatially decaying. As opposed to arc-cosine kernels, Matérn kernels are spatially decaying; they tend to zero as the distance $\|x-y\|$ becomes large. Consequently, although Matérn kernels are technically not compact (or locally supported), because their value decays exponentially fast (porcu2023), they can be considered effectively compact. From a computational perspective, this is a very desirable property as it leads to sparse Gram matrices when truncated with a special kernel (genton2001), allowing the use of highly efficient and scalable sparse linear solvers. Finally, we note that simply truncating any kernel does not yield a valid, positive definite kernel in general.
@@ 3.2 Eigenvalue Decay, Spectral Bias, and Reconstruction Error
Eigenvalue decay and spectral bias. We begin by showing that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings help coordinate-based MLPs overcome spectral bias. As investigated in tancik2020, a rapid decrease of the NTK’s eigenvalues implies an associated MLP’s slow convergence to high-frequency components of the target function (up to the point where the network is practically unable to learn these components). Consequently, a smaller eigenvalue decay rate (EDR), i.e., a slower eigenvalue decay, leads to faster convergence to high-frequency content—in the context of implicit surface reconstruction, more detailed geometry. To overcome this slow convergence, a phenomenon known as spectral bias, tancik2020 use a Fourier feature mapping of the form
@@
Theorem 1 (seeger2008, Theorem 3).
@@
Figure 3 visualizes the EDR for different values of $\nu$ and $h$ along with the EDR of the non-tunable, first-order arc-cosine kernel which equates to $\Theta(s^{-(1+d)/d})$ and has only been very recently proven by li2024 for general input domains and distributions. As seen, the smoother the kernel (i.e., the larger $\nu$) and the greater $h$, the faster the eigenvalues decay. This implies less detailed surface reconstruction (with a higher error) for smoother kernels and larger $h$; indeed, this is exactly what we observe in practice, see Figure 4. Contrary, by lowering both parameters, one can achieve faster convergence to high-frequency geometric details, effectively leading to more nuanced surface reconstruction and a smaller reconstruction error.
Figure 4: Matérn kernels are tunable. Surface reconstructions can be improved in practice by varying the bandwidth parameter, $h$, effectively tuning the kernels’ EDR (see also Figure 3). However, setting $h$ too small (<0.5<0.5) results in overfitting, while setting $h$ too big (>10>10) oversmooths (i.e., underfits) the true surface. This is also reflected in the reconstruction error (measured using Chamfer distance), plotted as a function of $h$ on the right.
@@
is the norm of the RKHS induced by a Matérn kernel with bandwidth $h$ and smoothness $\nu$. Moreover, $C_{d,\nu}:=(2^{d}\pi^{d/2}\Gamma(\nu+d/2)(2\nu)^{\nu})/\Gamma(\nu)$, and $\mathcal{F}[f]$ denotes the Fourier transform of a function $f$. For more information on the bound, including details on $C_{\mathcal{X},\Omega}$, please see, e.g., santin2016. By inspecting Eq. (10), we notice that the magnitude of the RKHS norm has a considerably high influence on the reconstruction error. Based on this observation and the fact that $\|f\|_{H_{\nu}}$ depends on $h$, our goal is now to further study the effect of $h$ on the RKHS norm. As a first result, we have:
@@ 3.3 Relation to Neural Networks and the Arc-cosine Kernel
Relation to neural networks. Next, we study the connection between Matérn kernels and neural networks. It is well known that the first-order arc-cosine kernel mimics the computation of a two-layer, infinite-width ReLU network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$, $f(x)=m^{-1/2}\sum_{i=1}^{m}v_{i}[w^{\top}_{i}x+b_{i}]_{+}$ when the bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$ are fixed from initialization and drawn from a standard normal distribution, see, e.g., cho2009; williams2021. Here, $x]_{+}:=\max\{0,x\}$ denotes the ReLU activation function.
Under similar assumptions, we now show that Matérn kernels mimic the computation of two-layer Sirens (sitzmann2020) if the layer-width approaches infinity. We claim the following:
@@
Theorem 4 (chen2021, Theorem 1).
@@
The RKHS of the Laplace kernel is equal to the RKHS induced by a sum of zeroth and first-order arc-cosine kernels, implying that, when restricted to $\mathbb{S}^{d-1}$, the Laplace and first-order arc-cosine kernel are equal up to constant scaling and addition of another kernel function.
See Appendix E for a proof. For general input domains, our empirical results presented next support (at least for ($L=2$)-layer networks) the widely believed claim that the NTK is not significantly different from standard kernels (belkin2018; hui2019; geifman2020), such as the Laplace kernel. Indeed, we show that the Laplace kernel even outperforms the arc-cosine kernel.